Calculate the number of atoms of carbon (C) in 463 cm3 of the colorless gas cyclobutane at 0 °C and atmospheric pressure, where its density is 2.50×10-3 g cm-3. The molecular formula of cyclobutane is C4H8.
Answer #1

Volume = 463 cm3 = 0.463 L (1cm3 = 0.001 L)

Molar mass of C4H8 = 4C + 8H = 4(12) + 8(1)

= 48+8 = 56 g/mol

Now, at STP (0°C, 1atm) :

22.4 L contain 1 mole C4H8 = 6.022×1023 molecules

0.463 L contain 0.463×6.022×1023/22.4 molecules

= 1.24×1022 molecules

1 molecule of C4H8 contain 4 C atoms

1.24×1022 molecule of C4H8 contain 4×1.24×1022 atoms

= 4.98×1022 C atoms